your cousin is 8 years older than your brother. three years ago, your cousin was twice as old as your brother. how old is your cousin now? how old is your brother now?
Answer by lwsshak3(11628)

your cousin is 8 years older than your brother. three years ago, your cousin was twice as old as your brother. how old is your cousin now? how old is your brother now?
***
let x=brother's age
x+8=cousin's age
..
3 yrs ago:
x-3=brother's age
x+8-3=x+5=cousin's age
x+5=2(x-3)
x+5=2x-6
x=11
x+8=19
brother's age=11
cousin's age=19